One Watt is (Select all that apply)

A

1 kg m2 / s3

B

1 J / s  

C

1 kg m / s3       

D

1 kg m / s2   

E

1 N m / s
Answer #1

A )

1 Watt = 1 kg * m^2 / s^3

J Nm kg m2 3

B)

1 Watt = J /s

J = Kg * m^2 / s^2 = N * m = W * s

J = W* s

W = J /s

E )

1 Watt = N m /s

the option c and d is not correct